OpenStudy (anonymous):

Find a polynomial function that has given zeros, 0, -2, -3 I understand the -2 and -3, but when I solve for the zero is it just x-0?

OpenStudy (jdoe0001):

well x = -2 => x +2 = 0 => (x+2) = 0 x = -3 => x + 3 = 0 => (x+3) = 0 x = 0 => x +0 = 0 =>(x) = 0

OpenStudy (jdoe0001):

0 is neither negative nor positive, so it doesn't matter what sign you use

OpenStudy (anonymous):

And then you simply foil?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Thanks! And one more quick thing, how do u solve when the zeros are 1+sqrt3 and 1-sqrt3

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.